a clyinderical hall has an smooth horizantal floor.a ball is projected along the floor from a point on the wall in an directiom making an angle x with the radius through the point.the ball return 2 the inital points after two impacts with the wallif the coefficent of restution is e,then tan^2x is
ans is e^3/e^2+e+1
how can i get this ans

Grade:

1 Answers

Badiuddin askIITians.ismu Expert
14 years ago

Dear jimmyu

let after impact angel made by ball with the radius is y

and after second impact angel made by ball with the radius is z

so 2x+2y+2z =180

or  x+y+z = 90 ........................1

at the first point where ball made first impact

 vsinx = v1siny

and  e = v1cosy /vcosx

 

so e = tanx/tay    or  tany = tanx/e  ...........2

similerly for second impact

      tanz = tany/e  or   tanz = tanx /e2 ............3

 

from equation 1

(90-x) = y+z

or tan(90-x) = tany + tanz /(1-tanytanz)

  1/tanx = (tany + tanz)/(1-tanytanz)

 

put value of  tany and tanz from equation 2and 3

tan2x = e3/(e2 +e +1)

Please feel free to post as many doubts on our discussion forum as you can.
If you find any question Difficult to understand - post it here and we will get you
the answer and detailed solution very quickly.

 We are all IITians and here to help you in your IIT JEE preparation.

 All the best.
 
Regards,
Askiitians Experts
Badiuddin